Solve the equation by first clearing the fraction or decimals.


43. -3/7c = -6
Any help with this equation would be Greatly Appreciated.

Answer by checkley71(8403)

-3C/7=-6
-3C=-42
C=-42/-3
C=14 ANSWER.